OpenStudy (anonymous):

Factor the expression by taking out the GCF. 3x - 9x3 ,my "answer" 3(x - 3x3) either that or it starts with an x and then the parentheses

OpenStudy (thomas5267):

What is 9x3? This \(9x^3\) or \(9x(3)=27x\)

OpenStudy (anonymous):

\[ 3x(1−3x^2)=3x(1−x\sqrt {3}\, )(1+x \sqrt {3} ) \]